You borrow a $ 345,000 add-on interest loan from the credit union and will repay in equal installments over 12 years. The nominal rate of interest is 4.75 %. Assuming quarterly repayment and simple rate of interest, obtain the equal quarterly payments.
Answer #1

Total interest to be paid = principal*interest rate*term of loan

= 345,000*4.75%*12 = 196,650

Total payment to be made = principal + interest = 345,000 + 196,650 = 541,650

Number of payments to be made = payments per year*term = 4*12 = 48

Equal quarterly payment = Total payment/Number of payments = 541,650/48 = 11,284.38 (Answer)